OpenStudy (anonymous):

What is the solution to the equation? -8x + 9x - 21= -12-16 x= ? I got 7 on this one too

OpenStudy (agl202):

Simplifying -8x + 9x + -21 = -12 + -16 Reorder the terms: -21 + -8x + 9x = -12 + -16 Combine like terms: -8x + 9x = 1x -21 + 1x = -12 + -16 Combine like terms: -12 + -16 = -28 -21 + 1x = -28 Solving -21 + 1x = -28 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '21' to each side of the equation. -21 + 21 + 1x = -28 + 21 Combine like terms: -21 + 21 = 0 0 + 1x = -28 + 21 1x = -28 + 21 Combine like terms: -28 + 21 = -7 1x = -7 Divide each side by '1'. x = -7 Simplifying x = -7 Is ur answer!

OpenStudy (anonymous):

Thank You again :)

OpenStudy (agl202):

Ur welcome!